OpenStudy (anonymous):

Is there a way to solve this with one variable? The sum of Carl's age and Ellen's age is 40 years. Carl's age 10 years from now will be 1 year less than 4 times Ellen's age 6 years ago. Find their present ages.

OpenStudy (anonymous):

Yes, try expressing Carl's age in terms of Ellen's and tell me what you get.

OpenStudy (anonymous):

Let the present ages of Carl be "c" and Ellen "e". thus we can develop this: c + e = 40 implying c = 40 - e

OpenStudy (anonymous):

Thanks for the help. The wording of the chapter was first-degree equations in one variable. Got the answer now.

OpenStudy (anonymous):

x = ellen age now 40 - x = carl age now 50 - x = 4(x - 6) - 1 50 - x = 4x - 25 75 = 5x x = 15 ellen = 15 carl = 25 thanks again